# Waveform Optimization and Beam Focusing for Near-field Wireless Power Transfer with Dynamic Metasurface Antennas and Non-linear Energy Harvesters

### v2

@@ V Numerical Analysis
In this section, we provide numerical analysis of the system performance. We consider an indoor office with a transmitter located at the center of the ceiling. The operating frequency of the system is $f_{1}=5.18$ GHz, which matches the characteristics of the utilized rectifier model [36]. The spacing between the elements in the DMA is $\lambda_{1}/5$, while $\lambda_{1}/2$ is the distance between two consecutive microstrips. Meanwhile, the inter-element distance is $\lambda_{1}/2$ in the fully-digital architecture. Thus, $N_{v}=N_{h}=\lfloor\frac{L}{\lambda_{1}/2}\rfloor$ for the fully-digital system, and $N_{v}=\lfloor\frac{L}{\lambda_{1}/2}\rfloor,N_{h}=\lfloor\frac{L}{\lambda_{1}/5}\rfloor$ for the DMA-assisted system [27]. Note that $L$ is the array length while the arrays are considered to be square-shaped. We set the optimization parameters $\tau_{s}=10^{-3}$, $\varsigma=5$, and $\upsilon=10^{-6}$. Without loss of generality, we set $G=1$ and $\bar{\eta}=\frac{\pi}{4}$ [^5] In practice, the HPA output power is larger than the input power because $G>1$. However, the proposed framework applies to all values of $G$.. Finally, the rectifier parameters are $v_{t}=25$ mV and $\eta_{0}=1.05$ [36, 21, 20].
@@
Fig. 5 provides the convergence performance of the proposed ASCA-DMA approach by presenting the amount of power consumption at the end of each iteration of alternating optimization. It is seen that the objective value gradually decreases with iterations until convergence, while the number of required iterations for convergence depends on the setup. For instance, it is shown that increasing $L$, $N_{f}$, or $M$ can increase the complexity of the problem leading to more iterations. Fig. 6 provides a detailed algorithm performance by showing all iterations of optimization, including the SCA algorithm for both optimizing metamaterials and digital precoders. It is seen that since the minimum harvested power increases when optimizing $\mathbf{q}$, the power consumption increases, while this facilitates decreasing the amount of power consumption when optimizing digital precoders. Therefore, the power consumption decreases gradually at the end of each iteration of the alternating optimization (red arrows) until convergence.
Figure 5: The convergence performance of ASCA-DMA for (a) $M=1$, $L=15$ cm, $N_{f}=1$ (left), (b) $M=1$, $L=25$ cm, $N_{f}=1$ (middle), and (c) $M=2$, $L=25$ cm, $N_{f}=4$ (right) over iterations with $d=2.5$ m and $P_{max}=100$ W.
Figure 6: The convergence performance of ASCA-DMA for $M=1$, $L=25$ cm, $d=2.5$ m, $N_{f}=1$, and $P_{max}=1$ W over all iterations (alternating and SCA). The red and green arrows indicate the starting points of optimization for the metamaterial elements and digital precoders, respectively.
Fig. 7 showcases the power consumption of the system as a function of the antenna length. Note that increasing the $L$ reduces the power consumption since the number of elements and the array aperture increases. This leads to more capability in beam focusing, thus, delivering more power to the devices given the same transmit power. However, the performance comparison between DMA and FD highly depends on the system setup. It is seen that DMA outperforms FD when $L$ and $P_{max}$ are relatively low. Note that the output of each RF chain in DMA has to feed all the corresponding elements, while the the amount of power consumption scales with the saturation power of the HPA. Thus, when $P_{max}$ is low, the output power of each RF chain in DMA is multiplied by a small value, leading to potential performance gains compared to FD. However, this is only the case when $L$ is low and the number of RF chains is small. On the other hand, in FD, $N_{rf}$ increases with $L$ at a higher rate compared to DMA. Thus, it is easier to reduce the amount of output power of the HPAs by distributing the required output power among them, leading to lower power consumption compared to DMA. However, the value of $L$ that shifts the favorable architecture from DMA to FD or vice versa depends on $P_{max}$ and the transmit power required for meeting the EH requirements. For instance, when $d=2.5$ m and $P_{max}=1$ W, the required transmit power and the power consumption multiplier are both low, leading to a shift in performance with lower $L$. Meanwhile, when $d=6.5$ m and $P_{max}=100$ W, DMA outperforms FD up to a higher value of $L$ compared to the previous case. Meanwhile, when $d$ is small and $P_{max}$ is sufficiently large, FD outperforms DMA over all $L$ values with the performance gap increasing with $L$. In contrast, when $d$ is relatively large and $P_{max}$ is small, DMA becomes the favorable choice for sufficiently large $L$.
Figure 7: The power consumption as a function of $L$ for $N_{f}=4$, $P_{max}\in\{1,100\}$ W, and $d\in\{2.5,6.5\}$ m.
Our results in Fig. 8 corroborate that increasing $N_{f}$ reduces $P_{c}$. As discussed in Section II-C, this is because the HPAs operate in the linear regime, and higher $N_{f}$ can leverage the rectifier non-linearity and deliver more DC power to the ERs via waveform optimization. However, as previously mentioned, the preference for DMA or FD highly depends on the system parameters. For instance, when $L=10$ cm and $P_{max}=1$ W, both $N_{rf}$ and saturation power have small values, leading to a lower power consumption for DMA compared to FD. The reason is that the multiplier of the HPA output power is low, thus, a lower value is multiplied by the output power of each RF chain in DMA. Combining this with a small $L$, thus a small $N_{rf}$ for FD, leads to DMA outperforming FD. On the other hand, different parameters and system setups may lead to different performances. For example, it is seen that when $L$ and saturation power are both sufficiently large, the favorable architecture shifts from DMA to FD. The reason is that the number of RF chains is much higher for FD in this case, leading to lower output power for each. Then, combining this with a large value of saturation voltage leads to smaller HPA outputs in FD, thus, lower power consumption. Meanwhile, increasing $N_{f}$ affects the performance gap between FD and DMA, and increasing $N_{f}$ leads to more performance gains in FD compared to DMA since the number of signals is relatively larger in FD. Thus, when DMA outperforms FD for a given $L$ and $P_{max}$, FD may start outperforming DMA by sufficiently increasing $N_{f}$. On the other hand, when FD performs better, increasing $N_{f}$ can increase the performance gap.
Figure 8: The power consumption as a function of $N_{f}$ for $L=25$ cm, $P_{max}\in\{1,100\}$ W, $L\in\{10,20\}$ cm, and $d=2.5$ m.
Fig. 9 illustrates the impact of user distance on performance. It is obvious that the power consumption increases with distance since the path loss becomes larger and more transmit power is required to overcome that and meet the EH requirements. Interestingly, the performance shift between FD and DMA is also shown here. It is shown that when both $P_{max}$ and $L$ are relatively low, DMA outperforms FD, especially over large distances. On the other hand, when $L$ increases, DMA starts performing better after a certain $d$ since the number of HPAs is much larger in FD, and increasing their output power affects the power consumption considerably. For instance, when $L=20$ cm and $P_{max}=100$ W, FD starts with better performance than DMA, but the performance gap decreases as $d$ becomes larger.
Figure 9: The power consumption as a function of user distance for $N_{f}=4$, $P_{max}\in\{1,100\}$ W, and $L\in\{10,20\}$ cm.
The impact of the number of users on system performance has been illustrated in Fig. 10 for different system parameters. As expected, the power consumption increases with $M$ since more EH requirements must be met. As previously shown, both DMA or FD may be the favorable choice depending on the system setup and the number of devices. For example, when $P_{max}$ and $L$ are both sufficiently large, DMA outperforms FD for a small $M$, but FD becomes favorable when $M$ is relatively large. The reason is that increasing $M$ leads to more required transmit power to satisfy the requirements and as earlier mentioned, a large $L$ leads to relatively smaller output powers compared in FD to DMA.
[…]
Figure 10: The power consumption over $M$ for (a) $P_{max}=1$ W (top) and (b) $P_{max}=100$ W (bottom), while $N_{f}=4$, and $L\in\{15,25\}$ cm. The users are located at $d=4.4$ m.
Fig. 11 provides some insights regarding the beam focusing capability in the near-field WPT by illustrating the normalized received RF power in each spatial point of the area. Note that the received signal at each point is normalized by its path loss to remove the impact of the distance. In Fig. 11a, it is seen that when the device is located in the near-field region, the beam is focused around the device location, while the beam trace fades increasingly past the device. Such phenomena can have a huge benefit in reducing the RF emission footprint in the environment, which facilitates the implementation of environmentally friendly WPT systems. Meanwhile, the beam pattern for a located device in the far-field is formed in the user’s direction, as illustrated in Fig. 11b. This may be highly disadvantaged in interference-sensitive applications since the generated beam may cause difficult-to-handle interference in the signals conveying information, e.g., in SWIPT.
Figure 11: The normalized received RF power (W) in the area when the energy receiver is located at (a) the near-field region (top) and (b) the far-field region (bottom) in the DMA-assisted system with $L=30$ cm and $N_{f}=1$.
The received RF waveform in DMA-assisted and fully-digital systems is presented in Fig.12a and Fig.12b, respectively. As previously mentioned, high PAPR waveforms are beneficial for enhancing the performance in terms of DC harvested power [7]. Moreover, our simulations verify this by showing that the received signal experiences high peak amplitudes at specific intervals. Recall that when HPAs are operating in the linear regime, as in our case, the HPA does not introduce distortion to the signal. Thus, it is beneficial to utilize multiple tones to leverage the rectifier’s non-linearity. Note that the peak-to-peak time depends on the characteristics of the EH circuit, mainly the capacitor.
Figure 12: The received signal at the device using (a) DMA (top) and (b) FD (bottom) for $M=1$, $L=25$ cm, and $N_{f}=8$.
VI Conclusion and Future Work
In this paper, we investigated a multi-antenna near-field WPT system with a DMA as the transmitter to charge multiple non-linear EH devices. Furthermore, we proposed an optimization framework relying on alternating optimization and SCA for the joint waveform optimization and beam focusing to minimize the system power consumption while meeting the EH requirements. Numerical results showed that both DMA and fully-digital architecture may be the favorable choice in terms of power consumption depending on the system setups and parameters such as antenna length, saturation power of the HPAs, number of users, and user distance. Moreover, we showed that increasing the antenna length or the number of tones can enhance the performance. Finally, we verified that the transmitter can focus the energy beams on the spatial points in the near-field region, while energy beams are formed toward the devices’ direction in the far-field.
As a prospect for future research, we may delve deeper into the signal generation aspect by analyzing the power consumption based on the number of tones. Another research direction is to utilize optimization approaches with lower complexity, e.g., relying on machine learning, to learn online the input-output relation of the system’s non-linear components while optimizing the transmit waveform accordingly.
-A Proof of Theorem 2
We relax (32b) by limiting the values of $q_{i,l}$ to lie within the Lorentzian circle in the complex plane. By utilizing the fact that (32a) is a positive and increasing function of the rectifier’s output voltage and using the epigraph form, the relaxed problem can be written as (2). Note that each configuration of a metamaterial corresponds to a point on the Lorentzian-constrained circle. Imagine $\vec{e}$ is the vector that represents the direction and gain of a point on the Lorentzian circle, while this direction and gain impacts the transmit signal. Meanwhile, the goal of (2) is to increase the minimum output voltage of the ERs. Therefore, when shaping the transmit signal toward different ERs, it is obvious that $\vec{e}$ should be chosen with the most possible gain in the required direction to improve the signal strength at the receiver. Furthermore, the most possible introduced gain by a metamaterial element along a specified direction happens when the point is exactly on the Lorentzian-constrained circle in that direction. Hence, although (33b) is a relaxed version of the constraint (32b), considering the final solution of the metamaterials as $a\vec{e},0\leq a\leq 1$, the only solution that leads to the most gain for the desired direction is $a=1$. Thus, the solution of (2) is the same as (III-B2) leading to a configuration positioned on the Lorentzian circle and the equivalence between the problems.

### v3

@@ V Numerical Analysis
In this section, we provide numerical analysis of the system performance. We consider an indoor office with a transmitter located at the center of the ceiling. The operating frequency of the system is $f_{1}=5.18$ GHz, which matches the characteristics of the utilized rectifier model [39]. The spacing between the elements in the DMA is $\lambda_{1}/5$, while $\lambda_{1}/2$ is the distance between two consecutive microstrips. Meanwhile, the inter-element distance is $\lambda_{1}/2$ in the fully-digital architecture. Thus, $N_{v}=N_{h}=\lfloor{2L}/{\lambda_{1}}\rfloor$ for the fully-digital system, and $N_{v}=\lfloor{2L}/{\lambda_{1}}\rfloor,N_{h}=\lfloor{5L}/{\lambda_{1}}\rfloor$ for the DMA-assisted system [30]. Note that $L$ is the array length while the arrays are considered to be square-shaped. We set the optimization parameters $\tau_{s}=10^{-3}$, $\varsigma=5$, and $\upsilon=10^{-6}$. Without loss of generality, we set $G=1$ and $\bar{\eta}=\frac{\pi}{4}$ [^6] In practice, the HPA output power is larger than the input power because $G>1$. However, the proposed framework applies to any values of $G$.. Finally, the rectifier parameters are $v_{t}=25$ mV and $\eta_{0}=1.05$ [39, 24, 23].
We utilize the characteristics of the ®Rogers RO4000 series ceramic laminate to calculate the propagation coefficients of the microstrips. Specifically, we calculate the attenuation and propagation coefficients of a RO400C LoPro with a thickness of 20.7 mil (0.5258 mm) using the formulation provided in [26], which gives $\alpha=0.356$ m-1 and $\beta=202.19$ m-1. Based on the rectifier circuit design and simulations in [39], the rectifier circuit diode enters the breakdown region when the received RF power is approximately $100\ \mu$W for a continuous wave ($N_{f}=1$). Moreover, it has been shown that the maximum RF-to-DC conversion efficiency is approximately $20\%$ for the mentioned setup. Hence, we establish a minimum requirement of $\tilde{P}_{dc}=20\ \mu$W for DC harvested power. In the figures, FD refers to the fully-digital architecture. Moreover, $d$ represents the distance between the ER and the center of the transmitter. Note that since the utilized problem formulation in (19) is equivalent to maximizing the PTE, the showcased results for power consumption inherently capture the PTE of the system. Specifically, the PTE and power consumption are inversely proportional.
In order to showcase the power consumption ($P_{c}$) in the figures, we first get the solution, i.e., the metamaterials’ frequency response and the signal weights, by leveraging the proposed optimization framework. Then, $P_{c}$ is computed by sampling (18) using the Nyquist sampling theorem. Specifically, we consider a bandwidth $\textrm{BW}=10$ MHz and sub-carrier spacing $\Delta f=\textrm{BW}/N_{f}$. Then, the sampling rate is computed based on the maximum frequency, such that $f_{s}=2f_{N_{f}}$ is the sampling rate, $\Delta t=1/f_{s}$ is the sampling interval, and the signal duration is $1$ ms.
Fig. 5 provides the convergence performance of the proposed ASCA-DMA approach by presenting the power consumption at the end of each iteration of alternating optimization. It is seen that the objective value gradually decreases with the iterations until convergence, while the number of required iterations for convergence depends on the setup. For instance, it is shown that increasing $L$, $N_{f}$, or $M$ can increase the complexity of the problem, leading to more iterations. Fig. 6 provides a detailed algorithm performance by showing the results and the end of each iteration of optimization (both alternating and SCA), including the SCA algorithm for both optimizing metamaterials and digital precoders. It is seen that since the minimum harvested power increases when optimizing $\mathbf{q}$, the power consumption increases, while this facilitates decreasing the power consumption when optimizing the digital precoders. Therefore, the power consumption decreases gradually at the end of each iteration of the alternating optimization (red arrows) until convergence.
Figure 5: The convergence performance of ASCA-DMA for different $L$, $M$, and $N_{f}$ over iterations with $d=2.2$ m and $P_{max}=1$ W.
Figure 6: The convergence performance of ASCA-DMA for $M=1$, $L=25$ cm, $d=2.2$ m, $N_{f}=1$, and $P_{max}=1$ W over all iterations (alternating and SCA). The red and green arrows indicate the starting points of optimization for the metamaterial elements’ frequency response and digital precoders, respectively.
Fig. 7 showcases the power consumption of the system as a function of the antenna length. Note that increasing $L$ reduces the power consumption since the number of elements and the array aperture increases. This leads to a better beam focusing capability, thus, delivering more power to the ERs given the same transmit power. Although DMA outperforms FD, the performance gap between DMA and FD highly depends on the system setup. For instance, the gap is larger for higher $P_{max}$ due to the larger number of RF chains and HPAs in FD. Moreover, observe that the performance gains caused by increasing $L$ depend also on the distance of the receiver from the transmitter. Specifically, when the distance becomes larger, the power consumption decreases more substantially by utilizing larger antenna arrays.
Figure 7: The power consumption as a function of $L$ for $N_{f}=8$, $P_{max}\in\{1,100\}$ W, and $d\in\{2.2,4\}$ m.
Our results in Fig. 8 corroborate that increasing the number of transmit tones reduces the power consumption. As discussed in Section II-C, this is because the HPAs operate in the linear regime, and higher $N_{f}$ can leverage the rectifier non-linearity and deliver more DC power to the ERs via waveform optimization. However, as previously mentioned, the DMA architecture requires less power consumption, but the performance gap between DMA and FD highly depends on the system parameters. Meanwhile, observe that increasing $N_{f}$ leads to more performance gains in FD compared to DMA since the number of signals/RF chains is relatively larger in FD, and this provides additional degrees of freedom in designing the transmit signal. Thus, the performance gap between DMA and FD becomes smaller as $N_{f}$ increases.
Figure 8: The power consumption as a function of $N_{f}$ for $P_{max}\in\{1,100\}$ W, $L\in\{10,20\}$ cm, and $d=2.2$ m.
Fig. 9 illustrates the impact of the distance $d$ of the ER from the transmitter on the system performance. It is obvious that the power consumption increases with the distance since the path loss becomes larger and more transmit power is required to overcome it and meet the EH requirements. Interestingly, it can be seen that the performance gains caused by increasing the antenna length are more substantial when the receiver is located at larger distances.
Figure 9: The power consumption as a function of ER distance for $N_{f}=8$, $P_{max}\in\{1,100\}$ W, and $L\in\{10,20\}$ cm.
The impact of the number of ERs on system performance is illustrated in Fig. 10 for different system parameters. As expected, the power consumption increases with $M$ since more EH requirements must be met, leading to more required transmit power. As previously shown, DMA outperforms FD, but the performance gap between these architectures varies with the system setup and parameters such as antenna length and HPAs’ saturation power.
Figure 10: The power consumption over $M$ for (a) $P_{max}=1$ W (top) and (b) $P_{max}=100$ W (bottom), while $N_{f}=4$, and $L\in\{10,20\}$ cm. The ERs are located at $d=3.1$ m.
Fig. 11 provides some insights regarding the beam focusing capability in the near-field WPT by illustrating the normalized received RF power in each spatial point of the area. Note that the received signal at each point is normalized by its path loss to remove the impact of the distance. In Fig. 11a, it is seen that when the ER is located in the near-field region, the beam is focused around its location, while the beam trace fades increasingly past the ER. Such phenomena can have a huge benefit in reducing the RF emission footprint in the environment, which facilitates the implementation of environmentally friendly WPT systems [3]. Meanwhile, the beam pattern for a located device in the far-field is formed in the receiver’s direction, as illustrated in Fig. 11b. This may be highly disadvantaged in interference-sensitive applications since the generated beam may cause difficult-to-handle interference in the signals conveying information, e.g., in SWIPT.
Figure 11: The normalized received RF power (W) in the area when the energy receiver is located at (a) the near-field region (left) and (b) the far-field region (right) in the DMA-assisted system with $L=30$ cm and $N_{f}=1$.
Fig. 12 compares the performance of the proposed initialization method with a random initialization approach. For the latter, the variables are generated randomly 1000 times, and the set of variables that leads to the least power consumption while satisfying the EH requirements is chosen as the initial point of the optimization. Moreover, if the EH constraints are not met within 1000 realizations, the random generation is continued until the EH requirements are met. It is seen that the initialization method proposed in Section IV is effective and can lead the optimization framework to provide significantly better results in both transmit architectures.
Figure 12: The power consumption using different initialization methods in DMA and fully-digital architectures as a function of $L$ for $M=1$, $P_{max}=1$ W, $d=2.2$ m, and $N_{f}=8$.
The received RF waveform for both transmit architectures is presented in Fig.13.a. As previously mentioned, high PAPR waveforms are beneficial for enhancing the performance in terms of DC harvested power [7], and our simulations verify this by showing that the received signal experiences high peak amplitudes at specific intervals. Note that the peak-to-peak time depends on the characteristics of the EH circuit, mainly the capacitor. Recall that when HPAs are operating in the linear regime, as in our case, the HPA does not introduce distortion to the signal. Thus, it is beneficial to utilize multiple tones to leverage the rectifier’s non-linearity. Moreover, the average amplitude of the frequency components among the RF chains and the sub-carrier is showcased in Fig.13.b. It is seen that all frequency components are utilized and power is allocated to them with different ratios. Moreover, observe that the transmit power is distributed among all RF chains in both architectures to efficiently leverage the degrees of freedom in focusing the signal toward the receiver and meeting the EH requirement with less power consumption.
Figure 13: (a) The received signal with DMA (top-left) and FD (top-right), and (b) the average of the amplitude of the frequency components of the RF chains output signals over the RF chains (bottom-left) and over the sub-carriers (bottom-right) for $M=1$, $L=20$ cm, and $N_{f}=8$.
Some discussions on the practical challenges and complexities of implementing both architectures are in order. Notice that many more RF chains are needed in an FD array compared to a DMA with the same size, as illustrated in Fig. 14. For instance, 64 RF chains are needed in a 25 $\times$ 25 cm2 FD array and just 9 for a DMA with the same size at $f_{1}=5.18$ GHz. Notably, as $L$ substantially increases, the number of required RF chains for an FD array increases drastically as well, making DMA the reasonable implementation option. However, this must be taken with caution as studies considering mutual coupling are necessary. Such practical considerations may impact the performance of the DMA, especially when the inter-element distance becomes smaller. On the other hand, both architectures might lose some performance gains due to the sub-optimality of the algorithm. Thus, the actual performance gaps and trade-offs might change, and obtaining optimal solutions for this problem is challenging and still an open topic for research.
Figure 14: The total number of (a) antenna elements (left) and (b) RF chains (right) as a function of antenna length $L$.
VI Conclusion and Future Work
In this paper, we investigated a multi-antenna near-field WPT system with a DMA-assisted transmitter to charge multiple non-linear EH devices. Furthermore, we proposed an optimization framework relying on alternating optimization and SCA for the joint waveform optimization and beam focusing to minimize the system power consumption while meeting ERs’ EH requirements. Numerical results showed that DMA-assisted architecture outperforms the fully-digital structure in terms of power consumption, while their performance gap depends on the system setups and parameters such as antenna length, saturation power of the HPAs, number of ERs, and ER distance to the transmitter. Moreover, we showed that increasing the antenna length or the number of tones can enhance the performance. Finally, we verified that the transmitter can focus the energy beams on the spatial points in the near-field region, while energy beams are formed toward the devices’ direction in the far-field.
As a prospect for future research, we may delve deeper into the signal generation aspect by analyzing the power consumption based on the number of tones. Another research direction is to utilize optimization approaches with lower complexity, e.g., relying on machine learning, to learn online the input-output relation of the system’s non-linear components while optimizing the transmit waveform accordingly. Moreover, we may consider a mutual coupling-aware optimization to investigate the impact of the coupling between the antenna elements in the DMA architecture, which requires novel solutions and might change the system performance depending on the inter-element distance. Additionally, the proposed frameworks can be extended to traditional hybrid and analog architectures consisting of phase shifters to investigate the power consumption of those architectures compared to fully-digital and DMA-assisted structures. Finally, proposing novel solutions for DMA-assisted SWIPT systems with non-linear energy harvesters can be another interesting research opportunity.
-A Proof of Theorem 1
We can write
$\sum_{i=1}^{N_{rf}}\mathbb{E}_{t}\biggl{\{}\sqrt{\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}}\biggr{\}}\stackrel{{\scriptstyle\textnormal{(a)}}}{{\mathstrut{\leq}}}\sum_{i=1}^{N_{rf}}\sqrt{\mathbb{E}_{t}\biggl{\{}\sum_{l=1}^{N_{h}}x^{DMA}_{i,l}(t)^{2}\biggr{\}}}$
=$\stackrel{{\scriptstyle\textnormal{(b)}}}{{\mathstrut{=}}}\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}\mathbb{E}_{t}\biggl{\{}x^{DMA}_{i,l}(t)^{2}\biggr{\}}}\stackrel{{\scriptstyle\textnormal{(c)}}}{{\mathstrut{=}}}\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}\sum_{n=1}^{N_{f}}\frac{G^{2}}{2}|\omega_{i,n}q_{i,l}h_{i,l}|^{2}},$ (39)
where (-A) comes from the Jensen inequality [40], (39) from the linearity of the mathematical average operator, and (39) from the equivalence between the average power of the signal in time and frequency domain. Furthermore, we can write
$\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}\sum_{n=1}^{N_{f}}|\omega_{i,n}q_{i,l}h_{i,l}|^{2}}\!=\!\sum_{i=1}^{N_{rf}}\sqrt{\sum_{l=1}^{N_{h}}|q_{i,l}h_{i,l}|^{2}}\sqrt{\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}},$ (40)
where $\sqrt{\sum_{n=1}^{N_{f}}|\omega_{i,n}|^{2}}$ represents the $l_{2}$ norm operator for the vector $[\omega_{i,1},\ldots,\omega_{i,N_{rf}}]^{T}$, which is convex, leading to the convexity of the upper bound for fixed $q_{i,l}$.
-B Proof of Theorem 2